OpenStudy (sen_turtle):

How would you factor m^2+m+1/4? Please show steps.

OpenStudy (daniel.ohearn1):

(m^2) + (m) + (1/4) is what you want factored?

OpenStudy (daniel.ohearn1):

The terms in the function are the first of all.

OpenStudy (daniel.ohearn1):

That would be different from (m^2)+((m+1)/4) or ((m^2)+(m+1))/4

OpenStudy (sen_turtle):

(m^2)+(m)+(1/4)

OpenStudy (daniel.ohearn1):

(m+2)(m-1) would give you your second m when multiplied, then all you need is to make turn the -2 into 1/4

OpenStudy (daniel.ohearn1):

1st account for your m^2 and m terms

OpenStudy (daniel.ohearn1):

Actually alternatively you could even factor it evenly into 2 factors. m must be in both factors to get m^2 you know. Then they other numbers will be add up to 1 and multiply together to get (1/4). Can you think of those 2 numbers?

OpenStudy (sen_turtle):

Would that answer be 1/2?

OpenStudy (daniel.ohearn1):

Yes

OpenStudy (daniel.ohearn1):

Acknowledging the formula for factoring (a^2)+(2ab)+(b^2)=(a+b)(a+b) In this situation 2ab= 1m and a^2 = 1m^2 therefore 2b=1m b=1/2 since a=m=1